OpenStudy (anonymous):

PLEASE HELP - Will give medal! (2x^2+3x-5)+(x^2+11x-7)=180 what is the value of x?

OpenStudy (dangerousjesse):

Solve for x over the real numbers: 3 x^2+14 x-12 = 180 Write the quadratic equation in standard form. Divide both sides by 3: x^2+(14 x)/3-4 = 60 Solve the quadratic equation by completing the square. Add 4 to both sides: x^2+(14 x)/3 = 64 Take one half of the coefficient of x and square it, then add it to both sides. Add 49/9 to both sides: x^2+(14 x)/3+49/9 = 625/9 Factor the left hand side. Write the left hand side as a square: (x+7/3)^2 = 625/9 Eliminate the exponent on the left hand side. Take the square root of both sides: x+7/3 = 25/3 or x+7/3 = -25/3 Look at the first equation: Solve for x. Subtract 7/3 from both sides: x = 6 or x+7/3 = -25/3 Look at the second equation: Solve for x. Subtract 7/3 from both sides: Answer: | | x = 6 or x = -32/3

OpenStudy (anonymous):

Thanks @DangerousJesse!